10g of H2O2 impure sample when reacted with KI (solution) produced 0.5g iodine.What is percentage purity of H2O2?(AI=127,AK=39)

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Write the balanced eqn. first,
H2O2   +   2 KI    ------------------>       I2   +  K2O   +    H2O
moles of I2 = 0.5/254 =  0.00196
from mole analysis,
moles of H2O2 would be: 0.00196
mass of H2O2 = 0.00196*34 = 0.066929 gm.
% purity =  0.066929/10 * 100 = 0.66929 %